OpenStudy (anonymous):

Given four functions, place them in order of their y-intercept, from highest to lowest.

OpenStudy (anonymous):

a. g(x), h(x), j(x), f(x) b. h(x), j(x), g(x), f(x) c. f(x), g(x), j(x), h(x) d. j(x), g(x), h(x), f(x)

OpenStudy (jack1):

hi @badsuns

OpenStudy (anonymous):

i think i'm understanding it now, thanks!

OpenStudy (jack1):

1 mistake though: g(x) = 3(20)^x but if x =0 g(x) = 3(20)^0 g(x) = 3(1) g(x) = 3

OpenStudy (anonymous):

oh okay, that makes a lot more sense. thanks!